OpenStudy (anonymous):

How do you write the number 4.3 × 10-^2 in standard form? 0.00043 4.3 0.043 0.43

OpenStudy (anonymous):

@CoconutJJ

OpenStudy (anonymous):

@CoconutJJ plz help its due tommorow i cant risk failing

OpenStudy (anonymous):

@bloopman can you help?

OpenStudy (coconutjj):

what grade are you in?

OpenStudy (anonymous):

8th and i rreaally need to pass plz help its due in a few hours

OpenStudy (anonymous):

How do you write the standard number 144,800,000,000 in scientific notation?

OpenStudy (coconutjj):

0.043

OpenStudy (anonymous):

1.448 × 1011 14.48 × 1011 1.448 × 1012 14.48 × 1012

OpenStudy (anonymous):

?????????????????? are xfor multiply

OpenStudy (anonymous):

plz help

OpenStudy (coconutjj):

what r u suppose to do ?

OpenStudy (anonymous):

How do you write the standard number 144,800,000,000 in scientific notation?

OpenStudy (anonymous):

Look ill give you medals everytime i see your on just plz help

OpenStudy (coconutjj):

1.448 x 10^11

OpenStudy (anonymous):

How do you write the number 4.933 × 10^9 in standard form? 493,300,000 493,300,000,000 4,933,000,000 4,933,000

OpenStudy (coconutjj):

3rd one

OpenStudy (anonymous):

How do you write the standard number 0.0098 in scientific notation? 98 × 10-^2 9.8 × 10-^5 9.8 × 10-^3 9.8 × 10-^2

OpenStudy (coconutjj):

3

OpenStudy (anonymous):

again ?????? are x

OpenStudy (coconutjj):

huh?

OpenStudy (anonymous):

How do you write the following expanded number in standard notation? 900 + 30 + 5 + 0.5 + 0.08

OpenStudy (anonymous):

@CoconutJJ

OpenStudy (coconutjj):

935.58

OpenStudy (anonymous):

How do you write the following standard number in expanded notation? 60,040.1

OpenStudy (coconutjj):

6000 + 40 + 0.1

OpenStudy (anonymous):

Convert the number below from exponential to standard notation. 5(10^4) + 6(10^3) + 9(10^2) + 9(10^1) + 5(10^0)

OpenStudy (coconutjj):

56995

OpenStudy (anonymous):

The speed of light is 186,000 miles per second. What is this speed in scientific notation?

OpenStudy (anonymous):

@CoconutJJ

OpenStudy (coconutjj):

okay srsly, I did that question already....

OpenStudy (anonymous):

@CoconutJJ i dont remember

OpenStudy (coconutjj):

well look back, i'm starting to think this is some sort of joke

OpenStudy (anonymous):

plz dont think tht

OpenStudy (anonymous):

@CoconutJJ you didnt answer it

OpenStudy (anonymous):

@CoconutJJ i went back and looked but u had left

OpenStudy (anonymous):

You @bloopman your freaking me out by staying here

OpenStudy (anonymous):

10^(-2) = 1/(10^2) = 1/100 multiply 1/100 by the floating point

OpenStudy (anonymous):

The speed of light is 186,000 miles per second. What is this speed in scientific notation?

OpenStudy (anonymous):

Can you help with this its a decimal

OpenStudy (anonymous):

@CoconutJJ

OpenStudy (anonymous):

@Outlaw

OpenStudy (anonymous):

@bloopman

OpenStudy (anonymous):

The speed of light is 186,000 miles per second. What is this speed in scientific notation?

OpenStudy (anonymous):

plz help with the 186,000 one jj i am seriously in trouble if i dont getthis done i have 2 hrs @CoconutJJ